OpenStudy (anonymous):

If the value of a discriminant is 13, what does that tell us about the type and number of roots? 1 rational root 2 rational roots 2 complex conjugate roots 2 irrational roots

OpenStudy (debbieg):

What do you know about the discriminant?

OpenStudy (uri):

@ladyb What do you think?

OpenStudy (debbieg):

Your solutions are given by: \[\Large x=\frac{ -b\pm \sqrt{b^2-4ac} }{ 2a}\] And the discriminant is the \(\Large b^2-4ac\) part of that, right?

OpenStudy (debbieg):

So think about what kind of number you will get from the quadratic equation, if the \(\Large b^2-4ac\) part is =13.

OpenStudy (debbieg):

In fact, think about what WOULD have to be true of the discriminant, for each of these cases: 1 rational root? 2 rational roots? 2 complex conjugate roots? 2 irrational roots? In each case, you can describe under what conditions on the discriminant, you will get the type of roots described. :)

OpenStudy (anonymous):

okay @ Debbie G im unsure how to solve this , can you walk me through the steps to solve it

OpenStudy (debbieg):

Do you know what the discriminant is? Do you know what the quadratic formula is? (I gave it above, it's the x= part). Tell me what that I posted above, you are unclear on, and we can go from there.

OpenStudy (anonymous):

no i dont no the quadratic formula @ Debbie G

OpenStudy (debbieg):

Well, you are obviously studying quadratic equations in this class (what is the class?), given that you are asked this problem. Does this look familiar? \[\Large x=\frac{ -b\pm \sqrt{b^2-4ac} }{ 2a}\]That is the equation for the roots - also called the "zeros" - of a quadratic equation, which is an equation of the form\[\Large ax^2+bx+c=0\] Does any of that ring a bell?

OpenStudy (debbieg):

In other words, when you have an equation that looks like this: \(\Large ax^2+bx+c=0\) And you want to SOLVE it (that is, find the values for x that make it true), you plug the a, b, and c into this: \[\Large x=\frac{ -b\pm \sqrt{b^2-4ac} }{ 2a}\] And that gives you the x's that solve it. You might get 1 rational solution, 2 rational solutions, 2 irrational solutions, or 2 complex solutions, depending on the stuff that is under the radical sign, the \(b^2-4ac\), which is called the discriminant.

OpenStudy (anonymous):

yes because i remember a,b, and c , can have any value except that a cant be 0.

OpenStudy (debbieg):

Right!! Good! The only reason we don't want a=0 is because then our \(x^2\) term would go away, and it isn't a quadratic equation anymore, it's just a plain ol' linear equation (which is perfectly fine, but a different type of equation.) :)

OpenStudy (debbieg):

Now when you look at \(\Large x=\dfrac{ -b\pm \sqrt{b^2-4ac} }{ 2a}\) keep in mind, you can also write that as \(\Large x=\dfrac{ -b }{ 2a} \pm\dfrac{ \sqrt{b^2-4ac} }{ 2a}\)

OpenStudy (anonymous):

okay i get that part but now what i do now that i know the Quadratic formula

OpenStudy (debbieg):

hold one sec....

OpenStudy (anonymous):

ok

OpenStudy (debbieg):

ok, so you will get different KINDS of numbers from that, depending on what kind of number \(b^2-4ac\) is. For example, if \(b^2-4ac\) =0, then it reduces to just the single solution, \(\dfrac{-b}{2a}\) , which is a rational number (it has to be, since a, b, and c are all integers). So I've just given you a hint: when \(b^2-4ac\) =0 there is one rational root. Does that make sense?

OpenStudy (anonymous):

no i still dont get it

OpenStudy (debbieg):

if \(\Large b^2-4ac=0\) then \(\Large x=\dfrac{ -b }{ 2a} \pm\dfrac{ \sqrt{0} }{ 2a}=\dfrac{ -b }{ 2a} \) which is ONE RATIONAL NUMBER. It has to be rational because a and b are integers. It is only ONE solution because the 0 discriminant means that the \(\pm\) part of the quadratic formula goes away. So \(\Large \text{discriminant}=b^2-4ac=0\) guarantees ONE RATIONAL SOLUTION. Tell me if that makes sense, and if not, what don't you understand before we go on?

OpenStudy (anonymous):

okay now i get it

OpenStudy (debbieg):

Good. The whole problem, or just what I wrote above? Do you want me to continue?

OpenStudy (anonymous):

i just understand what u wrote above

OpenStudy (debbieg):

Now looking at \(b^2-4ac\), under what circumstances - what KIND of values for \(b^2-4ac\), would \(\Large x=\dfrac{ -b }{ 2a} \pm\dfrac{ \sqrt{b^2-4ac} }{ 2a}\) give you: 2 rational roots? 2 complex conjugate roots? 2 irrational roots? What would have to be true of \(b^2-4ac\)? We see why when \(b^2-4ac=0\), we only got ONE root, so no need to worry about that case. For any other (non-zero) value of \(b^2-4ac\), we will certainly get 2 results, because of the \(\pm\). So all we really need to worry about then is, what would be true of \(b^2-4ac\) in order that \(\Large \dfrac{ \sqrt{b^2-4ac} }{ 2a}\) would be rational? would be complex? would be irrational?

OpenStudy (debbieg):

Start with complex. That's probably the easiest one. :) What must be true of \(b^2-4ac\) to make that expression above COMPLEX?

OpenStudy (anonymous):

i am un sure :(

OpenStudy (debbieg):

Do you know what a complex number is?

OpenStudy (anonymous):

A complex number is a number that can be expressed in the form a + bi, where a and b are real numbers and i is the imaginary unit, where i2 = −1.

OpenStudy (debbieg):

ok, good, you know how to google. ;) lol.... now, technically, ALL REAL numbers are also complex numbers (but not all complex numbers are real numbers), but for our purposes, I'm just saying "complex number" to mean a number that has an "i" in it. So how does that "i" get there? It gets there when we try to take a square root of a NEGATIVE number, right? \(\sqrt{-9}=i\sqrt{9}=3i \)

OpenStudy (debbieg):

So what has to be true about \(\sqrt{b^2-4ac}\) in order that \(\Large \dfrac{ \sqrt{b^2-4ac} }{ 2a}\) would be COMPLEX?

OpenStudy (anonymous):

you would have to make 2a 2i right?

OpenStudy (debbieg):

no....

OpenStudy (anonymous):

oh im confuse

OpenStudy (debbieg):

The 2a is 2a. It is, what it is. You get a complex number when you take a SQUARE ROOT of a NEGATIVE NUMBER. So \(\Large \dfrac{ \sqrt{b^2-4ac} }{ 2a}\) will be complex, if \(\Large b^2-4ac\) IS WHAT??

OpenStudy (anonymous):

if it were negative then it would be b^2 - 4aci , correct or am i way off ?

OpenStudy (debbieg):

well, you're getting there. :) If \(\Large b^2-4ac<0\) then \(\Large \sqrt{b^2-4ac}\) is a complex number, because you have something NEGATIVE under that square root. So that means that \(\Large x=\dfrac{ -b }{ 2a} \pm\dfrac{ \sqrt{b^2-4ac} }{ 2a}\) gives us 2 COMPLEX SOLUTIONS which are the "complex conjugates". So now you know when you get 1 rational solution, and when you get 2 complex conjugate solutions. And neither case happens when the discriminant, \(\Large b^2-4ac=13\). So what's left is, what does it take to get: 2 rational roots? 2 irrational roots?

OpenStudy (debbieg):

We know that it isn't \(\Large b^2-4ac\le0\) because we've covered those cases above. so we know \(\Large b^2-4ac>0\), which by the way, is true for \(\Large b^2-4ac=13\). So the only question is, what about \(\Large b^2-4ac\) makes the two solutions RATIONAL NUMBERS vs. IRRATIONAL NUMBERS?? Well, a rational number is just a number of the form \(\Large \dfrac { m }{n}\)where m and n are integers. So how would \(\Large \dfrac{ \sqrt{b^2-4ac} }{ 2a}\) be an integer? We already know that 2a is an integer, so we just need \(\Large \sqrt{b^2-4ac}\) to be an integer, which it will be if the radical part simplifies to an integer, which it will if the stuff under the square root (THE DISCRIMINANT) is a perfect square. If the stuff under the square root is NOT a perfect square, then you can't get rid of that square root, so you are left with an IRRATIONAL NUMBER. For example: \(\Large \sqrt{16}=4\) so if \(\Large b^2-4ac=16\) (or any perfect square) then you will have 2 rational roots. But \(\Large \sqrt{20}=\sqrt{4\cdot5}=2\sqrt{5}\), so if \(\Large b^2-4ac=20\) (or any positive number that is NOT a perfect square, like say 13) then you will have 2 IRRATIONAL roots.

OpenStudy (debbieg):

@ladyb I know I threw a lot of stuff at you in this thread. I could have just given you the answer 2 hours ago, but hopefully you learned something from our discussion and have gained a better understanding of quadratic functions and their roots. If so, you will have an easier time answering a question like this on your own, next time. :)

OpenStudy (anonymous):

i appreciate you for teaching me and not just giving me the answer i now have a better understanding of quadratic equations :)